### v1

# MiniConGTS: A Near Ultimate Minimalist Contrastive Grid Tagging Scheme for Aspect Sentiment Triplet Extraction

Abstract
Aspect Sentiment Triplet Extraction (ASTE) aims to co-extract the sentiment triplets in a given corpus. Existing approaches within the pretraining-finetuning paradigm tend to either meticulously craft complex tagging schemes and classification heads, or incorporate external semantic augmentation to enhance performance. In this study, we, for the first time, re-evaluate the redundancy in tagging schemes and the internal enhancement in pretrained representations. We propose a method to improve and utilize pretrained representations by integrating a minimalist tagging scheme and a novel token-level contrastive learning strategy. The proposed approach demonstrates comparable or superior performance compared to state-of-the-art techniques while featuring a more compact design and reduced computational overhead. Additionally, we are the first to formally evaluate GPT-4’s performance in few-shot learning and Chain-of-Thought scenarios for this task. The results demonstrate that the pretraining-finetuning paradigm remains highly effective even in the era of large language models.
1 Introduction
Aspect-Based Sentiment Analysis (ABSA) aims to jointly extract opinion terms, aspect terms (targets of the corresponding opinions), and their specific sentiment polarities in a given corpus. In the milestone research by Peng et al. (2020), the compound ABSA subtasks were consolidated into the Aspect Sentiment Triplet Extraction (ASTE) framework. For each input corpus, ASTE outputs triplets in the form (Aspect, Opinion, Polarity), where the Aspect term is the target or entity being discussed, the Opinion term is the sentiment or opinion expressed about the aspect, and Polarity indicates whether the opinion is positive, negative, or neutral. Figure 1 illustrates the ASTE task.
Figure 1: An illustration for ASTE, given the sentence "Bob Dylan is a great rocker, despite the broken CDs.", there are three triplets to be extracted: (Bob Dylan, great, positive), (rocker, great, positive), (CDs, broken, negative).
As an emerging fine-grained sentiment analysis initiative, ASTE offers a more detailed and nuanced understanding of sentiments in text compared to traditional methods that provide only an overall sentiment score Peng et al. (2020). This aspect-level structured approach is inherently more challenging.
Previous approaches to ASTE have generally followed two paradigms: Pipeline methods and Joint Tagging methods Zhang et al. (2022a). Pipeline methods decompose the ASTE task into multiple sequential subtasks, often suffering from error propagation Xu et al. (2020). Recent progress in Machine Reading Comprehension (MRC) also contributes to this paradigm Zhai et al. (2022); Mao et al. (2021); Zou et al. (2024); Chen et al. (2021b). Joint tagging methods adopt a unified tagging scheme to extract all triplet elements in one stage Xu et al. (2020). The key idea is to design a tagging scheme Zheng et al. (2017) that simultaneously predicts aspect terms, opinion terms, and sentiment polarities. Further developments have introduced a Grid Tagging Scheme (GTS) to represent the triplets on a unique 2D tableWu et al. (2020a); Zhang et al. (2022b); Chen et al. (2021b, 2022); Fei et al. (2022).
Recent advances in these approaches have been focusing on the classification head design Chen et al. (2022); Zhang et al. (2022b) and external semantic information enhancing Chen et al. (2021b, 2022); Fei et al. (2022); Jiang et al. (2023); Iswari et al. (2024). However, existing research has neglected the synergistic optimization within the joint tagging scheme and the integration of contextual word representations. In this research, we propose a method to effectively improve and utilize the representation capabilities of pretrained encoders in ABSA by integrating a minimalist tagging scheme and a novel token-level contrastive learning approach.
The proposed approach demonstrates comparable or superior performance in comparison to state-of-the-art techniques, while featuring a more compact design and reduced computational overhead. Notably, even in the era of Large Language Models (LLMs), our method exhibits superior effectiveness compared to GPT 3.5 and GPT 4 in both few-shot and Chain-of-Thought Wei et al. (2022) learning scenarios. This study provides valuable insights for the advancement of ASTE techniques within the paradigm of LLMs. Overall, our contributions are summarized as follows:
  1. 1.
    Minimalist Grid Tagging Scheme: We propose a novel minimalist joint tagging scheme that uses the fewest label classes to date.
  2. 2.
    Token-level Contrastive Learning Strategy: We introduce a token-level contrastive learning framework that enhances the contextual embeddings produced by the pretrained model. This framework is seamlessly geared towards our minimalist Grid Tagging Scheme (GTS) to effectively address the ASTE task.
  3. 3.
    Comprehensive Evaluation: We conduct extensive experiments and evaluations on multiple benchmark datasets, demonstrating the effectiveness and superiority of our proposed methods over existing approaches. Notably, we are the first to reveal GPT-4’s performance on this task, showcasing our method’s superior efficiency and effectiveness in the era of large language models.
2 Literature Review
2.1 ASTE Paradigms
Peng et al. (2020) proposed a pipeline method that divides ASTE tasks into two stages: initially extracting (Aspect, Opinion) pairs and subsequently predicting sentiment polarity. However, pipeline methods typically suffer from error propagation issues Xu et al. (2020). Recent pipeline methods treat ASTE as a Machine Reading Comprehension problem, and develops seq2seq methods such as machine reading comprehension Zhai et al. (2022); Mao et al. (2021); Zou et al. (2024); Chen et al. (2021b). Joint Tagging strategies are remarked by certain Unified Tagging Scheme designs, where elements of a triplet can be extracted simultaneously. ET Xu et al. (2020) introduced a position-awared tagging scheme with a conditional random field module, effectively addressing span overlapping issues. Recent joint paradigm methods have refined the ASTE task with the development of proficient Grid Tagging Schemes (GTS).
2.2 Grid Tagging Scheme
Wu et al. (2020a) pioneered the adoption of a grid tagging scheme (GTS) for ASTE, yielding substantial performance gains. Subsequent research refined and enhanced GTS. BDTF Zhang et al. (2022b) designed a boundary-driven tagging scheme, effectively reducing boundary prediction errors. Alternative research augmented GTS by integrating external semantic information as structured knowledge into their models. $\mathrm{S}^{3}\mathrm{E}^{2}$ Chen et al. (2021b) retained the GTS tagging scheme while introducing novel semantic and syntactic enhancement modules between word embedding outputs and the tagging scheme. EMGCN Chen et al. (2022) incorporated external knowledge from four areas—Part-of-Speech Combination, Syntactic Dependency Type, Tree-based Distance, and Relative Position Distance—through an exogenous hard-encoding strategy. SyMux Fei et al. (2022) contributed a unified tagging scheme capable of handling all ABSA subtasks by integrating insights from GCN, syntax encoders, and representation multiplexing.
2.3 Contrastive Learning
While contrastive learning has gained popularity in diverse NLP domains Wu et al. (2020b); Giorgi et al. (2021); Gao et al. (2021); Zhang et al. (2021), its application to ASTE remains relatively unexplored. Ye et al. (2021) adopts contrastive learning into triplet extraction in a generative fashion. Wang et al. (2022) takes contrastive learning as a data augmentation approach. Yang et al. (2023) proposed an enhancement approach in pairing with two separate encoders.
3 Method
3.1 Overall Framework
An overall description of the training process can be found in Figure 2. Basically, our design can be broken down into the Minimalist Grid Tagging Scheme (GTS) and the Token-level Contrastive Learning Strategy.
Tokenize the input sequence S using the Tokenizer $\mathrm{Tk}$ and pass the tokenized sequence through the Pretrained Language Model $\mathrm{PLM}$ (such as BERT) to obtain contextualized representations h:
$\textit{{h}}=\mathrm{PLM}\left(\mathrm{Tk}\left(\textit{{S}}\right)\right).$ (1)
Then, the inference phase involves with forming the Minimalist GTS and predicting the corresponding class for each cell. Once the GTS is predicted, it can be decoded by the GTS decoder into triplets in natural language form. The training phase also introduces a novel contrastive learning strategy, in which similar and dissimilar pairs of contextual representations are distinguished. The contrastive loss is then weighted and summed with the tagging loss, which is the classification loss between the predicted and ground-truth tagging schemes.
Our research benefits from the following two closely intertwined aspects: 1) The use of the Minimalist GTS simplifies the learning process by reducing the number of labels, facilitating faster convergence, and seamlessly gearing the contrastive learning. 2) The token-level contrastive learning enhances the model’s ability to distinguish between related and unrelated elements within the input sequence, thereby improving the overall accuracy of the tagging system. For a more detailed description of our algorithm pipeline, see the pseudo-code in Appendix A.1.
Figure 2: An overview of the proposed method, where the “Encoder” denotes for the sequential combination of a Tokenizer and a Pretrained Language Model (PLM).
3.2 Minimalist Grid Tagging Scheme
3.2.1 Tagging Scheme Design
As defined by Section 3.1, once an input sentence is encoded into a sequence of contextual representations $\textit{{h}}=\{h_{1},h_{2},...,h_{|\textit{{h}}|}\}$, we form a $|\textit{{h}}|\times|\textit{{h}}|$ matrix, that is, our tagging scheme $\textit{{tag}}_{|\textit{{h}}|\times|\textit{{h}}|}$. As shown in Figure 3, on the rows we mark Aspect tokens by yellow and the columns we mark Opinion tokens by green (positive) and blue (negative). Then the each intersection of these marked rows and columns can uniquely represent an identical sentiment triplet. Thus, each such triplet can be denoted by a 2-D area (submatirx) in the matrix, where Sentiment Polarity is indicated with POS. (positive), NEU. (neutral), or NEG. (negative) in the top-left corner cell of the area, while CTD. indicates the continuation of the pairing relationship within the same region. MSK. (mask) on the diagonal represents masked cells that are not involved in the computation. In Figure 3, an example sentence is tokenized and tagged.
By defining our grid tagging scheme, we frame the triplet extraction problem as a 5-class classification task, using the fewest number of labels known to date. In Appendices A.3 - A.5, we provide rigorous proof and heuristic insights to justify our design and ensure its rationality.
Figure 3: The grid tagging scheme employs the fewest classes of labels while completely handle all the triplet cases without conflict, overlap or omission. Each area circled in red dashed lines corresponds to a triplets. For example, intersection area between columns of "broken" and rows of "CDs" is marked as negative, with NEG. on its top-left cell and CTD. for others. It is worth mentioning that the blank cells in the matrix are labeled as an additional class but are omitted for visual simplicity.
3.2.2 Tagging Loss
We adopt a tagging loss to guide neural network learning. We concatenate the representation with its transposed form to construct a matrix. Then, we apply the classification head $\mathrm{Cls}$ to the embeddings, followed by the softmax function, to obtain the predicted classification probabilities for each cell:
$\hat{\textit{{tag}}}_{i,j}=\text{softmax}(\mathrm{Cls}(\textit{{h}}_{i};\textit{{h}}_{j}^{\mathrm{T}})),i,j=1,...,|\textit{{h}}|.$ (2)
The focal loss Lin et al. (2017) is employed to mitigate class imbalance by placing greater emphasis on examples that are difficult to classify correctly. This is achieved by down-weighting the loss for well-classified instances and focusing more on misclassified instances. The formula for focal loss $\mathcal{L}$ is as follows:
$\mathcal{L}_{\text{tag}}=-\frac{1}{|\textit{{h}}|^{2}}\sum_{i,j=1}^{|\textit{{h}}|}\alpha_{\textit{{tag}}_{i,j}}(1-\textit{tag}_{t})^{\gamma}\log(\textit{tag}_{t}),$ (3)
where $\alpha$ is a weighting factor for balancing the importance of tags, $\gamma$ is a focusing parameter that increases the weight of hard-to-predict tags, and $\textit{{tag}}_{i,j}$ and $tag_{t}$ represent the ground truth label and the predicted probability for the true label at position $(i,j)$, respectively:
$tag_{t}=\hat{\textit{{tag}}}_{i,j;\textit{{tag}}_{i,j}}$ (4)
3.3 Contrstive Learning Strategy
3.3.1 Contrastive Learning Label Matrix
Contrastive learning is an unsupervised learning method that aims to learn effective feature embeddings by pulling together similar pairs of samples and pushing apart dissimilar pairs. In our design, we construct a label matrix where each cell is annotated by either PULL or PUSH, which means making the representations closer among tokens within the same class and farther between those of different classes. See an illustration of this strategy in Figure 4.
Figure 4: An illustration for the “Contrastive Mask”. Each token is paired with every other token, where PULL denotes positive sample pairs, indicating that the tokens belong to the same category and should be pulled closer together, while PUSH denotes negative sample pairs, indicating that the tokens belong to different categories and should be pushed apart. The lower triangular part of the matrix, marked by MSK. are masked cells that are not involved in the computation. For example, "Bob" and "Dylan" are marked as a positive sample pair with PULL, indicating similarity, while "Bob" and "is" are marked as a negative sample pair with PUSH, indicating dissimilarity.
3.3.2 Objective Function
The commonly used InfoNCE (Information Noise-Contrastive Estimation) loss function van den Oord et al. (2019) is employed:
[…]
where $\textit{{h}}_{i}^{+}$ / $\textit{{h}}_{j}^{-}$ represents the positive / negative sample embedding with the anchor respectively. $\text{sim}(\cdot,\cdot)$ denotes the similarity function, which is calculated by the cosine similarity:
$\text{sim}(\textit{{h}}_{i},\textit{{h}}_{i})=\frac{\textit{{h}}_{i}\cdot\textit{{h}}_{i}}{\|\textit{{h}}_{i}\|\|\textit{{h}}_{i}\|}$ (6)
3.4 Overall Loss Function
The overall loss $\mathcal{L}$ can be formulated as a weighted sum of two individual loss functions: the tagging loss $\mathcal{L}_{tag}$ and the contrastive loss $\mathcal{L}_{contrast}$:
$\mathcal{L}=\mathcal{L}_{tag}+\beta\mathcal{L}_{contrast},$ (7)
where $\beta$ is a coefficient for balancing the different parts of the loss. This combination allows for balancing the influence of each component in the training process.
4 Experiments
4.1 Implementation Details
All experiments were performed on a single RTX 2080 Ti. The best model weight on the development set is saved and then evaluated on the test set. For the PLM encoder, the pretrained weights bert_base_uncased and roberta_base are downloaded from (Wolf et al., 2020). GPT 3.5-Turbo and GPT 4 are implemented using OpenAI API OpenAI (2024). The learning rate is $1\times 10^{-5}$ for the PLM encoder, and $1\times 10^{-3}$ for the classification head.
4.2 Datasets
We evaluate our method on two canonical ASTE datasets derived from the SemEval Challenges Pontiki et al. (2014, 2015, 2016). These datasets serve as benchmarks in most aspects-based sentiment analysis (ABSA) research. The first dataset, denoted as $\mathcal{D}_{1}$, is the Aspect-oriented Fine-grained Opinion Extraction (AFOE) dataset introduced by Wu et al. (2020a). The second dataset, denoted as $\mathcal{D}_{2}$, is a refined version by Xu et al. (2020), building upon the work of Peng et al. (2020). More details are provided in Appendix A.2.
4.3 Baselines
We evaluate our method against various techniques including pipeline, sequence-labeling, seq2seq, table-filling and LLM-based approaches. Detailed descriptions for each method can be found in the Appendix A.6.
4.4 ASTE Performance
4.4.1 Comparison to Existing Methods
Methods 14Res 14Lap 15Res 16Res
P R F1 P R F1 P R F1 P R F1
Pipeline
[…]
Sequence-tagging
Span-BART Yan et al. (2021) 65.52 64.99 65.25 61.41 56.19 58.69 59.14 59.38 59.26 66.60 68.68 67.62
JET Xu et al. (2020) 70.56 55.94 62.40 55.39 47.33 51.04 64.45 51.96 57.53 70.42 58.37 63.83
Seq2seq
Dual-MRC Mao et al. (2021) 71.55 69.14 70.32 57.39 53.88 55.58 63.78 51.87 57.21 68.60 66.24 67.40
$\textrm{BMRC}^{\dagger}$ Chen et al. (2021a) 72.17 65.43 68.64 65.91 52.15 58.18 62.48 55.55 58.79 69.87 65.68 67.35
COM-MRC Zhai et al. (2022) 75.46 68.91 72.01 62.35 58.16 60.17 68.35 61.24 64.53 71.55 71.59 71.57
Triple-MRC Zou et al. (2024) - - 72.45 - - 60.72 - - 62.86 - - 68.65
Table-filling
GTS Wu et al. (2020a) 67.76 67.29 67.50 57.82 51.32 54.36 62.59 57.94 60.15 66.08 66.91 67.93
Double-encoder Jing et al. (2021) 67.95 71.23 69.55 62.12 56.38 59.11 58.55 60.00 59.27 70.65 70.23 70.44
EMC-GCN Chen et al. (2022) 71.21 72.39 71.78 61.70 56.26 58.81 61.54 62.47 61.93 65.62 71.30 68.33
BDTF Zhang et al. (2022b) 75.53 73.24 74.35 68.94 55.97 61.74 68.76 63.71 66.12 71.44 73.13 72.27
STAGE-1D Liang et al. (2023) 79.54 68.47 73.58 71.48 53.97 61.49 72.05 58.23 64.37 78.38 69.10 73.45
STAGE-2D Liang et al. (2023) 78.51 69.3 73.61 70.56 55.16 61.88 72.33 58.93 64.94 77.67 68.44 72.75
STAGE-3D Liang et al. (2023) 78.58 69.58 73.76 71.98 53.86 61.58 73.63 57.9 64.79 76.67 70.12 73.24
DGCNAP Li et al. (2023) 72.90 68.69 70.72 62.02 53.79 57.57 62.23 60.21 61.19 69.75 69.44 69.58
LLM-based
GPT-3.5-turbo zero-shot 44.88 55.13 49.48 30.04 41.04 34.69 36.02 53.40 43.02 39.92 57.78 47.22
GPT-3.5-turbo few-shots 51.51 65.19 57.55 39.79 50.09 44.35 43.34 63.09 51.39 51.12 71.01 59.45
GPT-3.5-turbo CoT 48.47 59.05 53.24 30.48 40.30 34.71 39.51 56.70 46.57 44.03 63.81 52.10
GPT-3.5-turbo CoT+few-shots 49.41 59.15 53.85 33.78 42.33 37.57 39.02 56.08 46.02 46.49 66.93 54.86
GPT-4o zero-shot 32.99 38.13 35.37 17.81 22.55 19.90 27.85 37.73 32.05 32.17 43.00 36.80
GPT-4o few-shots 54.11 66.20 59.55 38.23 48.61 42.80 45.57 60.41 51.95 52.90 71.01 60.63
GPT-4o CoT 41.21 53.32 46.49 26.98 37.71 31.46 33.07 50.93 40.10 39.14 58.17 46.79
GPT-4o CoT+few-shots 46.81 59.86 52.54 29.71 40.85 34.40 35.08 53.81 42.47 41.53 61.09 49.45
Ours
MiniConGTS 76.1 75.08 75.59 66.82 60.68 63.61 66.50 63.86 65.15 75.52 74.14 74.83
Table 1: Experimental results on $\mathcal{D}_{2}$ Xu et al. (2020). The best results are highlighted in bold, while the second best results are underscored. The results with $\dagger$ are retrieved from Yu Bai Jian et al. (2021). The results with $\natural$ are retrieved from Xu et al. (2020). The results with $\sharp$ are retrieved from Peng et al. (2020). The results with $\ddagger$ are retrieved from Mao et al. (2021).
We evaluate ASTE performance using the widely accepted (Precision, Recall, F1) metrics. The results of the data set $\mathcal{D}_{2}$ are shown in Table 1, while the results of $\mathcal{D}_{1}$ are presented in Appendix Table 8. The best results are highlighted in bold, and the second-best results are underscored. Our proposed method consistently achieves state-of-the-art performance or ranks second in most evaluated cases.
In particular, on the data set $\mathcal{D}_{1}$, the proposed method achieves a substantial improvement of 3. 08% in the F1 score in the 14Lap subset. This improvement is particularly significant given that the highest score on this subset is the lowest among all datasets, showcasing our method’s effectiveness in handling challenging instances. Moreover, on the 14Res subset, our F1 score exceeds 76.00, which, to the best of our knowledge, is the highest reported performance. For the dataset $\mathcal{D}_{2}$, our method outperforms all state-of-the-art approaches by more than 1 percentage point on the 14Res, 14Lap, and 16Res subsets. Only on the 16Res subset does the BDTF method (Zhang et al., 2022b) achieve a slightly better performance.
4.4.2 Comparison to GPT
Our proposed method is based on the Pretrain-Finetuning paradigm, which is increasingly challenged by large language models (LLMs) Kojima et al. (2022); Wei et al. (2021). It is concerned about how the advancing capabilities of LLMs might impact the ASTE task.
When compared to advanced LLMs, the performance and computational efficiency of our method stand out. As shown in Tables 8, 1, and 9, even the state-of-the-art LLM, GPT-4, with its staggering number of parameters, does not achieve satisfactory results for ASTE, even with few-shot learning and Chain-of-Thought (CoT) Wei et al. (2022) enhancement. Additionally, using LLMs introduces significant computational overhead. For more information on experiment setting see Appendix A.9. For detailed results see Table 1, 12, Appendix A.8 and Appendix A.10. Note that, fine-tuning LLMs may offer some improvements, but it also risks catastrophic forgetting Shi et al. (2024) and is left for future work.
To our knowledge, this is the first formal study to evaluate GPT-4’s performance on these ASTE datasets, providing valuable insights for future research.
4.5 Performance on Other ABSA Tasks
Our method can also effectively handle other ABSA subtasks, including Aspect Extraction (AE), Opinion Extraction (OE), and Aspect Opinion Pair Extraction (AOPE). AE aims to extract all the (Aspect) terms, OE aims to extract all the Opinion terms, and AOPE aims to extract all the (Aspect, Opinion) pairs from raw text. The results for these tasks are presented in Appendix A.7, where our method consistently achieves best F1-scores across nearly all tasks.
Models $\mathcal{D}_{1}$ $\mathcal{D}_{2}$
14Res 14Lap 15Res 16Res 14Res 14Lap 15Res 16Res
MiniConGTS 76.00 64.07 65.43 71.80 75.59 63.61 65.15 74.83
w/o. RoBERTa 74.12 63.18 62.95 69.41 72.66 62.15 63.25 70.71
      $\Delta$F_1 -1.88 -0.89 -2.48 -2.39 -2.93 -1.46 -1.90 -4.12
      w/o. contr 72.61 61.94 58.14 68.16 71.72 61.49 58.11 68.03
      $\Delta$F_1 -3.39 -2.13 -7.29 -3.64 -3.87 -2.12 -7.04 -6.80
      w/o. tag 67.78 54.98 60.75 62.62 65.83 54.98 58.73 67.63
      $\Delta$F_1 -8.22 -9.09 -4.68 -9.18 -9.76 -8.63 -6.42 -7.20
Table 2: Ablation study on F1, where “w/o. RoBERTa” denotes “Replace RoBERTa with bert-base-uncased”, “w/o. contr” denotes without the contrastive learning mechanism, and “w/o. tag” denotes “replace our tagging scheme with a baseline”.
Method Num Tags Features Enhancing
GTS Wu et al. (2020a) 6 None
Double-encoder Jing et al. (2021) 9 None
EMC-GCN Chen et al. (2022) 10 4 Groups
BDTF Zhang et al. (2022b) $2\times 2\times 3$ None
STAGE Liang et al. (2023) $2\times 2\times 4$ None
DGCNAP Li et al. (2023) 6 POS-tagging
Ours $5$ None
Table 3: Tagging Scheme Comparison.
5 Analysis
5.1 Ablation Study
In this section, we conduct a series of ablation experiments to demonstrate the superiority of our method and eliminate potential confounding factors. Experiments were conducted on the $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ datasets, using F1 scores as the comparison metric.
Encoder. We replaced the RoBERTa encoder with BERT, resulting in a slight decrease in F1 scores on both datasets, although our method still outperformed most other approaches.
Contrastive Learning. We deactivated the contrastive mechanism in our method (denoted as “w/o. contr”) by setting the coefficient of the contrastive loss to 0. The results in Table 2 illustrate a significant F1-score decrease of $2.12\sim 7.29$% in both datasets.
Tagging Scheme. We substituted our proposed scheme with the conventional GTS tagging scheme Wu et al. (2020a), resulting in a substantial performance decline (Table 2) of $4.68\sim 9.18$%. This indicates that the contrastive learning methods, within our framework, is of strong reliance on an appropriate tagging scheme. This reinforces the effectiveness of our compact yet impactful tagging scheme.
5.2 Effect of Contrastive Learning
Figure 5: A plot of the hidden word representation based on the $\mathcal{D}_{1}$ 14Res dataset, where the dimension is reduced to 3. “Pretrained” refers to the representation output by official released model. We finetune the pretrained model with and without contrastive learning strategy respectively.
In Figure 5, an example is shown illustrating how contrastive learning improves representation. The right upper-row subplots show the representation outputs with contrastive learning, while the lower row subplots display that without contrastive learning. Principal Component Analysis (PCA) (Maćkiewicz and Ratajczak, 1993) is used to reduce the vector dimensions to three for visualization purposes. The distributions indicate that contrastive learning significantly enhances the representations, with similar classes of hidden word representations becoming more tightly clustered and dissimilar classes more distinct.
5.3 Efficiency Comparison
Model Memory Num Params Epoch Time Inf Time F1(%) Device
Span-ASTE Xu et al. (2021) 3.173 GB - 108s - 71.62 Tesla v100 32GB
BDTF Zhang et al. (2022b) 8.103 GB $>$0.18B 135s - 74.73 Tesla v100 32GB
GPT 3.5-Turbo OpenAI (2024) $>$80 GB 175B - 0.83s 49.48 OpenAI API
GPT 4 OpenAI (2024) $>$80 GB 1760B - 1.56s 35.37 OpenAI API
Ours 7.11GB 0.12B 10s 0.01s 76.00 2080 Ti 11GB
Table 4: An efficiency comparison, where is evaluated by Gao (2021) and later confirmed by OpenAI Wikipedia (2024), is estimated by Schreiner (2023), is cited from Zhang et al. (2022b), and is reported by Wikipedia (2024). Epoch Time refers to the training time per epoch on the training set.
We compared the computational efficiency of MiniConGTS with other approaches, including baseline ASTE methods and LLMs, on an ASTE task. Evaluation metrics such as memory usage, number of parameters, epoch/inference time, and F1 scores are recorded in Table 4. Our approach not only requires less memory usage for higher performance compared to traditional ASTE methods but also offers much faster runtime even using a relatively lower-cost GPU.
Table 3 provides another comparative analysis of the tagging schemes. Our method has a compact design with the fewest classes of labels. Moreover, compared with other SOTA baselines, our method does not rely on any additional linguistic information enhancement.
5.4 Case Study
A case analysis is presented in Table 12, where the proposed method demonstrates solid performance. Despite minor faults in missing the full terms, it exhibits a profound understanding of the case.
It is quite interesting to investigate the GPT error cases. The findings reveal that the performance of the GPT model is mixed - while it is able to identify more aspect-opinion (A-O) pairs than the ground truth annotations in some cases, this comes at the cost of a decreased precision. This suggests that the GPT model may be “over-interpreting” the input, making inferences that go beyond what is strictly supported by the text. Furthermore, the GPT model appears to be overly sensitive to the presence of adverbs (such as “very”, “a bit”, etc.) in the input. This sensitivity manifests itself in the model frequently adding or removing adverbs when extracting the opinion components, which further contributes to a decrease in the overall accuracy of the ABSA task.
These findings highlight the importance of developing ABSA models that can strike the right balance between extracting all relevant aspect-opinion pairs, while still maintaining a high degree of precision. The effective use of encoding appears to be a promising direction for achieving this balance and advancing the state-of-the-art in Aspect-Based Sentiment Analysis.
6 Conclusion
In this work, we have introduced an elegant and efficient framework for ASTE to achieve SOTA performance. Our approach is built upon two effective components: a new tagging scheme and a novel token-level contrastive learning implementation. The ablation study demonstrates the synergy between these components, reducing the need for complex model designs and external information enhancements.
7 Limitations
Our method is based on a 2D matrix tagging scheme, where the time complexity for decoding, given the input corpus length $N$, is $O(N^{2})$. This may be unacceptable when $N$ is too large. Additionally, although we have demonstrated our method on commonly used classic English datasets, it should be tested on more natural corpora and for its cross-language capability.
8 Ethics & Potential Risks Statement
In our experiments, we used widely accepted datasets focused on e-commerce reviews, which have a lower risk of offensive content. We scrutinized the data for biases against gender, race, and marginalized groups. Despite these precautions, our model might still generate potentially offensive sentiment assessments if used inappropriately, such as evaluating ethical or moral statements. We reserve the right to limit or modify the use of our technology to prevent misuse.
Appendix A Appendix
A.1 Pseudo-code for the training process.
See Algorithm 1.
Algorithm 1 .
Modules:
Input:
Raw sentences: $\mathcal{S}_{|\mathcal{S}|}$;
Ground truth triplets: $\mathcal{T}_{|\mathcal{T}|}^{gt}$ , where
$\mathcal{T}_{k}=(A_{k},O_{k},S_{k})$, $k\in\{1,2,...,|\mathcal{T}|\}$;
classes of contrasted labels: $\mathcal{C}$.
Output:
Predicted Triplets: $\mathcal{T}_{|\mathcal{T}|}^{pred}$;
[…]
Algorithm:
Repeat for $N$ epochs:
1:  Hidden word representation: $\mathcal{H}_{|\mathcal{H}|}=\texttt{PLMsEncoder}(\mathcal{S}_{|\mathcal{S}|})$;
2:  Tensor Operations: ||×$\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}=\texttt{expand}(\mathcal{H}_{|\mathcal{H}|})$, ||×$\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T}=\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}.\texttt{transpose}()$;
3:  Similarity matrix:
$\mathbf{Sim}_{|\mathcal{H}|\times|\mathcal{H}|}=-(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})\circ(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})$
where
$\mathbf{Sim}_{i,j}=-\left\lVert\mathcal{H}_{i}-\mathcal{H}_{j}\right\rVert^{2}$
and $\circ$ denotes the Hadamard product.
4:  Contrastive Mask matrix: $\mathbf{M}_{|\mathcal{H}|\times|\mathcal{H}|}$, where   $\mathbf{M}_{i,j}=1$ if $\mathcal{H}_{i},mathcal{H}_{j}\in\mathcal{C}_{p},p\in{1,2,3}$ else $-1$;
5:  Contrastive loss: $\mathcal{L}_{contrastive}=\newline \sum_{i=1}^{|\mathcal{H}|}\sum_{j=1}^{|\mathcal{H}|}\big{(}\mathbf{Sim_{|\mathcal{H}|\times|\mathcal{H}|}}\circ\mathbf{M_{|\mathcal{H}|\times|\mathcal{H}|}}\big{)}_{i,j};$
6:  Predicted tagging matrix: $\mathbf{Tag}_{|\mathcal{H}|\times|\mathcal{H}|}^{pred}=\texttt{ClsHead}(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|},\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})$;
7:  Focal loss: $\mathcal{L}_{focal}=\texttt{FocalLoss}(\mathbf{Tag}_{|\mathcal{H}|\times|\mathcal{H}|}^{pred},\mathbf{Tag}_{|\mathcal{H}|\times|\mathcal{H}|}^{gt})$;
8:  Weighted Loss: $\mathcal{L}=\mathcal{L}_{focal}+\alpha\mathcal{L}_{contrastive}$.
9:  Backward propagation.
Predicted triplets:
$\mathcal{T}_{|\mathcal{T}|}^{pred}=\texttt{TaggingDecoder}(\mathbf{Tag}_{|\mathcal{H}|\times|\mathcal{H}|}^{pred})$
Metric:
[…]
A.2 Descriptive Statistics of The Datasets
See Table 5.
Datasets #S #A #O #S1 #S2 #S3 #T
14Res $\mathcal{D}_{1}$ Train 1259 1008 849 1456 164 446 2066
Dev 315 358 321 352 44 93 489
Test 493 591 433 651 59 141 851
$\mathcal{D}_{2}$ Train 1266 986 844 1692 166 480 2338
Dev 310 396 307 404 54 119 577
Test 492 579 437 773 66 155 994
14Lap $\mathcal{D}_{1}$ Train 899 731 693 691 107 466 1264
Dev 225 303 237 173 42 118 333
Test 332 411 330 305 62 101 468
$\mathcal{D}_{2}$ Train 906 733 695 817 126 517 1460
Dev 219 268 237 169 36 141 346
Test 328 400 329 364 63 116 543
15Res $\mathcal{D}_{1}$ Train 603 585 485 668 24 179 871
Dev 151 182 161 156 8 41 205
Test 325 353 307 293 19 124 436
$\mathcal{D}_{2}$ Train 605 582 462 783 25 205 1013
Dev 148 191 183 185 11 53 249
Test 322 347 310 317 25 143 485
16Res $\mathcal{D}_{1}$ Train 863 775 602 890 43 280 1213
Dev 216 270 237 224 8 66 298
Test 328 342 282 360 25 72 457
$\mathcal{D}_{2}$ Train 857 759 623 1015 50 329 1394
Dev 210 251 221 252 11 76 339
Test 326 338 282 407 29 78 514
Table 5: Statistic information of our two experiment datasets: “#S”, “#T”, “#A”, and “#O” denote the numbers of “Sentences”, “Triplets”, “Aspects”, and “Opinions”; “#S1”, “#S2”, #S3” denote the numbers of sentiments “Positive”, “Neutral” and “Negative”, respectively.
A.3 Rethinking the GTS
Rethinking the 2D tagging scheme:
Lemma 1. Specific to the ASTE task, when we take it as a 2D-labeling problem, we are to 1) find a set of tagging strategies to establish a 1-1 map between each triplet and its corresponding tagging matrix. See the proof in Appendix Proof 1.
Lemma 2. In a 2D-tagging for ASTE, at least three basic goals must be met: 1) correctly identifying the (Aspect, Opinion) pairs, 2) correctly classifying the sentiment polarity of the pair based on the context, and 3) avoiding boundary errors, such as overlapping [^1] It occurs when one single word belongs to multiple classes in different triplets. , confusion [^2] It occurs when there is a lack of location restrictions so that multiple neighbored candidates can not be uniquely distinguished. , and conflict [^3] It occurs when one single word is composed of multiple tokens, and the predict gives predictions that are not aligned with the word span. . See the proof in Appendix Proof 2.
Theorem 1. From insight of the above lemmas, it can be concluded that using enough (that is, following the 1-1 map properties in Lemma 1, as well as avoiding the issues in Lemma 2) labels will make it a theoretically ensured tagging scheme.
Assumption 1. Ceteris paribus, for a specific classification neural network, the fewer the number of target categories, the easier it is for the network to learn. This is a empirical and heuristic assumption, for the reasonable consideration of Simplification of Decision Boundaries Hinton and Salakhutdinov (2006) and Enhancement of Training Efficiency (less parameters).
Combining Theorem 1 and Assumption 1, fewer yet enough labels can be heuristically better solution with theoretical guarantee.
With the above knowledge, our tagging scheme employs a full matrix (illustrated as Figure 6) so that rectangular occupations in its cells indicate respective triplets, where each of the rectangles’ row indices correspond to the relative Aspect term and the column indices correspond to the Opinion. Hereafter, this kind of labels can be taken as a set of “place holder”, which is obviously a 1-1 map meeting Lemma 1.
To further satisfy Lemma 2, we introduce another kind of labels, “sentiment & beginning tag”. This set of labels specializes in recognizing the top-left corner of a “shadowed” area. Meanwhile, it takes a value from the sentiment polarity, i.e. Positive, Neutral, Negative. This tagging is crucial to both identify the beginning of an triplet and label the sentiment polarity.
Figure 6 shows a comprehensive case of our tagging scheme, in which the left matrix is an appearance of our tagging scheme and can be decomposed into two separate components. The middle matrix is the first component, which takes only one tag to locate the up-left beginning of an area, and the second component simply predicts a binary classification to figure out the full area.
Note that, this design benefits the tagging scheme’s decode process. By scanning across the matrix, we only start an examination function when triggered by a beginning label like this, and then search by row and column until it meets any label except a “continued” (“CTD”), which satisfies Lemma 2.
Figure 6: Decomposition of the tagging scheme into two components: 1) a beginning mark matrix with sentiment labels; and 2) a placeholder matrix denoting regions of triplets with “1”s and default regions with “0”s. Remember that each row is taken as candidates for an Aspect and each column is taken as candidates for an Opinion. Naturally, each cell in the square matrix can be seen as an ordered pair for a unique candidate of <Aspect, Opinion>. When we simply sum the two components up, we have the left-hand tagging scheme in Figure 6, where the “Sentiment & Beginning Tag” is like a trigger (just like you click your mouse), and the “Place Holder” is like a “continued shift” (continue to hold and drag the mouse to the downright).
A.4 Proof 1:
{adjustwidth}
0.9em0pt Let:
  • $S$ be a sentence with $n$ tokens.
  • $M$ be an $n\times n$ tagging matrix for $S$, where each entry $M[i][j]$ can hold a label.
  • $T_{k}=(A_{k},O_{k},S_{k})$ be a sentiment triplet consisting of an aspect term $A_{k}$, an opinion term $O_{k}$, and a sentiment $S_{k}$.
Tagging Strategy
If $A_{k}$ starts at position $i$ and $O_{k}$ starts at position $j$, then $M[i][j]$ is tagged with a unique label $L_{k}$ that encodes $S_{k}$. This label $L_{k}$ uniquely identifies the triplet $T_{k}$, ensuring that no other entry $M[i^{\prime}][j^{\prime}]$ with $(i^{\prime},j^{\prime})\neq(i,j)$ carries the same label unless it refers to the same sentiment context.
Define $\text{Define }L_{k}=\text{"start of triplet"}T_{k}\text{ with sentiment }S_{k}$
Proof of One-to-One Mapping
  • Injectivity: Each $L_{k}$ uniquely identifies a triplet $T_{k}$. If $M[i][j]=M[i^{\prime}][j^{\prime}]=L_{k}$, then by definition, $(i,j)=(i^{\prime},j^{\prime})$ and $T_{k}$ is the same.
  • Surjectivity: Each triplet $T_{k}$ can be uniquely located and identified by its label $L_{k}$ in matrix $M$, where no two distinct triplets have the same label at the same matrix position.
Conclusion
The tagging scheme ensures that each sentiment triplet $T_{k}$ is uniquely mapped to a specific label in the matrix $M$, and each label in $M$ uniquely refers back to a specific triplet $T_{k}$. This guarantees a one-to-one correspondence between the triplets and their tagging matrix representations, fulfilling the conditions required by Lemma 1 for an effective and efficient ASTE process.
A.5 Proof 2:
{adjustwidth}
0.9em0pt For the ASTE task, considered as a 2D-labeling problem, it is necessary to ensure three fundamental goals are met:
Definitions
  • $S$ be a sentence with $n$ tokens.
  • $M$ be an $n\times n$ tagging matrix for $S$, where each entry $M[i][j]$ can hold a label indicating a component of a sentiment triplet.
  • $T_{k}=(A_{k},O_{k},S_{k})$ be a sentiment triplet consisting of an aspect term $A_{k}$, an opinion term $O_{k}$, and a sentiment $S_{k}$.
Goals
  1. 1.
    Correct Identification of Pairs: Ensure that each (Aspect, Opinion) pair is correctly identified in the tagging matrix $M$.
  2. 2.
    Classification of Sentiment Polarity: Accurately classify the sentiment polarity $S_{k}$ for each (Aspect, Opinion) pair.
  3. 3.
    Avoidance of Boundary Errors: Prevent boundary errors such as overlapping and confusion in the tagging matrix $M$.
Proof Using Contraposition
  1. 1.
    Assuming Incorrect Identification: Assume that some (Aspect, Opinion) pairs are incorrectly identified in $M$. This would mean that there exists at least one pair $(i,j)$ where $M[i][j]$ does not represent the actual relationship (Aspect, Opinion) in $S$. This misrepresentation leads to incorrect sentiment analysis results, which contradicts the requirement of the task to provide accurate sentiment analysis, thereby proving that our identification must be correct.
  2. 2.
    Assuming Incorrect Classification: Assume the sentiment polarity $S_{k}$ is incorrectly classified in $M$. This would imply that the sentiment associated with an (Aspect, Opinion) pair is wrong, leading to a sentiment analysis that does not reflect the true sentiment of the text. Given that the primary goal of ASTE is to accurately identify sentiments, this assumption leads to a contradiction, thereby establishing that our classification must be accurate.
  3. 3.
    Assuming Existence of Boundary Errors: Assume boundary errors such as overlaps or confusion occur in $M$. Such errors would prevent the clear identification and classification of sentiment triplets, leading to incorrect or ambiguous extraction outcomes. This would undermine the integrity and usability of the ASTE process, contradicting the task’s need for precise extraction mechanisms. Hence, we prove that boundary errors must be effectively managed.
Conclusion
The contraposition approach solidifies that the ASTE tagging strategy in a 2D labeling framework successfully achieves the correct identification of pairs, accurate classification of sentiment, and effective management of boundary errors, as any failure in these aspects leads to contradictions with the task requirements.
A.6 Baselines
See Table 6.
Methods Brief Introduction
Pipeline
OTE-MTL Zhang et al. (2020) It proposes a multi-task learning framework including two parts: aspect and opinion tagging, along with word-level sentiment dependency parsing. This approach simultaneously extracts aspect and opinion terms while parsing sentiment dependencies using a biaffine scorer. Additionally, it employs triplet decoding based on the aforementioned outputs during inference to facilitate triplet extraction.
Li-unified-R+PD Peng et al. (2020) It proposes an unified tagging scheme, Li-unified-R, to assist target boundary detection. Two stacked LSTMs are employed to complete aspect-based sentiment prediction and the sequence labeling.
CMLA+C-GCN Wang et al. (2017) It facilitates triplet extraction by modelling the interaction between the aspects and opinions.
Two-satge Peng et al. (2020) It decomposes triplet extraction to two stages: 1) predicting unified aspect-sentiment and opinion tags; and 2) pairing the two results from stage one.
RI-NANTE+ Dai and Song (2019) It adopts the same sentiment triplets extracting method as that of CMLA+, but it incorporates a novel LSTM-CRF mechanism and fusion rules to capture word dependencies within sentences.
Sequence-tagging
Span-BART Yan et al. (2021) It redefines triplet extraction within an end-to-end framework by utilizing a sequence composed of pointer and sentiment class indexes. This is achieved by leveraging the pretrained sequence-to-sequence model BART to address ASTE.
JET Xu et al. (2020) It extracts triplets jointly by designing a position-aware sequence-tagging scheme to extract the triplets and capturing the rich interactions among the elements.
Seq2seq
Dual-MRC Mao et al. (2021) It proposes a solution for ASTE by jointly training two BERT-MRC models with parameters sharing.
BMRC Chen et al. (2021a) It introduces a bidirectional MRC (BMRC) framework for ASTE, employing three query types: non-restrictive extraction queries, restrictive extraction queries, and sentiment classification queries. The framework synergistically leverages two directions, one for sequential recognition of aspect-opinion-sentiment and the other for sequential recognition of opinion-aspects-sentiment expressions.
Table-filling
GTS Wu et al. (2020a) It proposes a novel 2D tagging scheme to address ASTE in an end-to-end fashion only with one unified grid tagging task. It also devises an effective inference strategy on GTS that utilizes mutual indication between different opinion factors to achieve more accurate extraction.
Double-encoder Jing et al. (2021) It proposes a dual-encoder model that capitalizes on encoder sharing while emphasizing differences to enhance effectiveness. One of the encoders, referred to as the pair encoder, specifically concentrates on candidate aspect-opinion pair classification, while the original encoder retains its focus on sequence labeling.
$\mathrm{S}^{3}\mathrm{E}^{2}$ Chen et al. (2021b) It represents the semantic and syntactic relationships between word pairs, employs GNNs for encoding, and applies a more efficient inference strategy.
EMC-GCN Chen et al. (2022) It employs a biaffine attention module to embed ten types of relations within sentences, transforming the sentence into a multi-channel graph while incorporating various enhanced linguistic features to enhance performance. Additionally, the method introduces an effective strategy for refining word-pair representations, aiding in the determination of whether word pairs are a match or not.
LLM-based
zero-shot Performing aspect-based sentiment analysis using an LLM. The specific method involves inputting a prompted sentence and directly outputting the corresponding [A, O, S] triplets. An example of the text given to the LLM, with the prompt added, is as follows: "Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. You only need to provide the triplets, no additional explanations are required. The provided text: {sentence}"
few-shots Building upon the zero-shot method, a small number of examples from the training set are added to the prompted sentence: "Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. For example: input: {train sentence} output: {train triplets}, … (some other examples). You only need to provide the triplets, no additional explanations are required. The provided text: {sentence}"
Table 6: Baseline methods with brief introduction.
A.7 Performance on Other ABSA Tasks
See Table 7.
Methods 14Res 14Lap 15Res 16Res
AE OE AOPE AE OE AOPE AE OE AOPE AE OE AOPE
CMLA 81.22 83.07 48.95 78.68 77.95 44.10 76.03 74.67 44.60 74.20 72.20 50.00
RINANTE 81.34 83.33 46.29 77.13 75.34 29.70 73.38 75.40 35.40 72.82 70.45 30.70
Li-unified 81.62 85.26 55.34 78.54 77.55 52.56 74.65 74.25 56.85 73.36 73.87 53.75
GTS 83.82 85.04 75.53 79.52 78.61 65.67 78.22 79.31 67.53 75.80 76.38 74.62
Dual-MRC 86.60 86.22 77.68 80.44 79.90 63.37 75.08 77.52 64.97 76.87 77.90 75.71
MiniConGTS (Ours) 86.55 87.04 79.60 82.62 83.41 73.23 86.53 83.05 73.87 85.48 87.06 76.29
$\Delta$F1 -0.05 0.82 1.92 2.18 3.51 7.56 8.31 3.74 6.34 8.61 9.16 0.58
Table 7: F1-score performance on other ABSA tasks: AE, OE, and AOPE. The test is implemented on $\mathcal{D}_{1}$. Results of other models are retrieved from Fei et al. (2022).
A.8 ASTE Performance on ($\mathcal{D}_{1}$)
See Table 8.
Methods 14Res 14Lap 15Res 16Res
P R F1 P R F1 P R F1 P R F1
Pipeline
OTE-MTL Zhang et al. (2020) - - 45.05 - - 59.67 - - 48.97 - - 55.83
[…]
RI-NANTE+ Dai and Song (2019) 31.42 39.38 34.95 21.71 18.66 20.07 29.88 30.06 29.97 25.68 22.30 23.87
[…]
Two-satge Peng et al. (2020) 58.89 60.41 59.64 48.62 45.52 47.02 51.7 46.04 48.71 59.25 58.09 59.67
Sequence-tagging
Span-BART Yan et al. (2021) - - 72.46 - - 57.59 - - 60.10 - - 69.98
JET Xu et al. (2020) 67.97 60.32 63.92 58.47 43.67 50.00 58.35 51.43 54.67 64.77 61.29 62.98
MRC based
BMRC Chen et al. (2021a) 71.32 70.09 70.69 65.12 54.41 59.27 63.71 58.63 61.05 67.74 68.56 68.13
COM-MRC Zhai et al. (2022) 76.45 69.67 72.89 64.73 56.09 60.09 68.50 59.74 63.65 72.80 70.85 71.79
Table-filling
$\mathrm{S}^{3}\mathrm{E}^{2}$ Chen et al. (2021b) 69.08 64.55 66.74 59.43 46.23 52.01 61.06 56.44 58.66 71.08 63.13 66.87
GTS Wu et al. (2020a) 70.92 69.49 70.20 57.52 51.92 54.58 59.29 58.07 58.67 68.58 66.60 67.58
EMC-GCN Chen et al. (2022) 71.85 72.12 71.78 61.46 55.56 58.32 59.89 61.05 60.38 65.08 71.66 68.18
BDTF Zhang et al. (2022b) 76.71 74.01 75.33 68.30 55.10 60.99 66.95 65.05 65.97 73.43 73.64 73.51
DGCNAP Li et al. (2023) 71.83 68.77 70.26 66.46 54.34 58.74 62.03 57.18 59.49 69.39 72.20 70.77
LLM-based
GPT-3.5-turbo zero-shot 39.21 56.17 46.18 26.21 40.69 31.88 31.21 52.75 39.21 35.28 59.64 44.34
GPT-3.5-turbo few-shots 50.32 64.75 56.63 29.67 43.90 35.41 36.94 61.01 46.02 44.80 69.96 54.62
GPT-3.5-turbo CoT 40.78 57.93 47.86 28.37 43.25 34.27 35.17 57.11 43.53 40.32 65.79 50.00
GPT-3.5-turbo CoT+few-shots 44.97 57.81 50.59 28.31 43.04 34.15 35.71 58.49 44.35 43.72 66.45 52.74
Ours
MiniConGTS 75.87 76.12 76.00 67.45 61.01 64.07 66.84 64.08 65.43 69.38 74.40 71.80
Table 8: Experimental results on $\mathcal{D}_{1}$ Wu et al. (2020a). The best results are highlighted in bold, while the second best results are underscored. The results with $\dagger$ are retrieved from Yu Bai Jian et al. (2021). The results with $\natural$ are retrieved from Xu et al. (2020). The results with $\flat$ are retrieved from Wu et al. (2020a). The results with $\sharp$ are retrieved from Peng et al. (2020).
A.9 Details Settings of GPT Experients
See Table 9.
Sentence Ground Truth Predictions
GPT-3.5 zero-shot GPT-3.5 few-shots GPT-4 zero-shot GPT-4 few-shots
It is a cozy place to go with a couple of friends. [place, cozy, positive] [place, cozy, positive],
[friends, couple of, positive]
[place, cozy, positive] [place, cozy, positive]
[going with friends, a couple of, positive]
[place, cozy, positive]
This place has the best sushi in the city. [sushi, best, positive] [sushi, best, positive],
[place, best, positive], [city, best, positive]
[sushi, best, positive] [sushi, the best, positive],
[place, This, neutral],
[city, in the, neutral]
[sushi, best, positive]
Disappointingly, their wonderful Saketini has been taken off the bar menu. [Saketini, wonderful, positive],
[bar menu, Disappointingly, negative]
[Saketini, wonderful, positive],
[bar menu, taken off, negative]
[Saketini, taken off, negative] [Saketini, has been taken off the bar menu, negative],
[bar menu, wonderful, positive]
[Saketini, wonderful, positive],
[bar menu, taken off, negative]
I went in one day asking for a table for a group and was greeted by a very rude hostess. [hostess, rude, negative] [hostess, rude, negative] [hostess, rude, negative] [table for a group, asking, neutral],
[hostess, very rude, negative]
[table for a group, asking, neutral],
[hostess, very rude, negative]
But make sure you have enough room on your credit card as the bill will leave a big dent in your wallet. [bill, big, negative] [credit card, enough room, positive],
[bill, big dent, negative],
[wallet, big dent, negative]
[credit card, enough room, negative],
[bill, big dent, negative],
[wallet, big dent, negative]
[room on your credit card, enough, positive],
[bill, will leave a big dent in your wallet, negative]
[bill, big dent, negative]
Table 9: In summary, there are several challenges observed in the performance of GPT models concerning triplets. Firstly, there is a prominent issue of "hard" matching, where GPT models tend to introduce additional modifiers or adverbs in the opinion component, leading to a lack of exact correspondence. Secondly, during zero-shot inference, GPT models tend to generate multiple predicted triplets, resulting in decreased precision. This behavior particularly hampers the precision of the model’s predictions. Thirdly, inconsistencies arise in handling triplets involving structures such as [A, O1 and O2, S] and [A, O1, S], [A, O2, S]. This inconsistency is challenging to mitigate due to its dependence on annotation practices and conventions. Upon closer examination, the issues observed do not appear to be as pronounced as indicated by the evaluation metrics. Rather, they often manifest as cases where the general idea is correctly captured, but the precise format or phrasing does not align perfectly. Notably, the performance of GPT-4 deteriorates due to its occasional tendency to not merely "extract" fragments from sentences but to generate its own summarizations. Consequently, evaluating against triplets that originate solely from annotated sentences poses a challenge in achieving alignment. Furthermore, GPT-4 exhibits a proclivity for extracting longer sequences of words as aspects or opinions, while GPT-3.5 tends to produce shorter sequences that better conform to typical annotation scenarios.
Methods Prompts
zero-shot Suppose you are an expert of aspect-based sentiment analysis. Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. You only need to provide the triplets, no additional explanations are required. The provided text: {sentence}
few-shots Suppose you are an expert of aspect-based sentiment analysis. Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. For example:
input: The food is uniformly exceptional , with a very capable kitchen which will proudly whip up whatever you feel like eating , whether it ’s on the menu or not .
output: [’food’, ’exceptional’, ’positive’], [’kitchen’, ’capable’, ’positive’]
(generated from training set)
Now I will provide a new sentence, and you only need to provide the triplets [Aspect, Opinion, Sentiment] without any additional explanations. The provided sentence: {sentence}
CoT Suppose you are an expert of aspect-based sentiment analysis. Please analyze the given text for aspect-based sentiment analysis using the following steps:
Definitions:
- Aspect: An aspect is a specific part or feature of the entity being discussed. It is usually a noun or a noun phrase.
- Opinion: An opinion is a descriptive term or phrase that expresses a sentiment towards the aspect. It is usually an adjective or a descriptive phrase.
- Sentiment: The sentiment is the overall feeling expressed towards the aspect, categorized as positive, negative, or neutral.
Instructions:
1. Read the text and identify all aspects mentioned.
2. For each identified aspect, determine the opinion expressed and the sentiment (positive, negative, neutral).
3. Summarize the findings in the format [Aspect, Opinion, Sentiment]. Each triplet must contain an aspect, an opinion, and a sentiment.
4. If there is a one-to-many relationship between aspects and opinions, list multiple triplets.
5. Use all words from the original text to answer without any changes.
Example: (automatically generated by ChatGPT-4o)
Text: "The restaurant has a great ambiance, but the service is poor and the food is average."
Steps:
1. Identify aspects: ambiance, service, food.
2. Evaluate opinions and sentiments:
- ambiance: Opinion - great, Sentiment - positive
- service: Opinion - poor, Sentiment - negative
- food: Opinion - average, Sentiment - neutral
3. Summarize:
- [ambiance, great, positive]
- [service, poor, negative]
- [food, average, neutral]
Please analyze the following text: {sentence}
CoT+few-shots Suppose you are an expert of aspect-based sentiment analysis. Please analyze the given text for aspect-based sentiment analysis using the following steps:
Definitions:
- Aspect: An aspect is a specific part or feature of the entity being discussed. It is usually a noun or a noun phrase.
- Opinion: An opinion is a descriptive term or phrase that expresses a sentiment towards the aspect. It is usually an adjective or a descriptive phrase.
- Sentiment: The sentiment is the overall feeling expressed towards the aspect, categorized as positive, negative, or neutral.
Instructions:
1. Read the text and identify all aspects mentioned.
2. For each identified aspect, determine the opinion expressed and the sentiment (positive, negative, neutral).
3. Summarize the findings in the format [Aspect, Opinion, Sentiment]. Each triplet must contain an aspect, an opinion, and a sentiment.
4. If there is a one-to-many relationship between aspects and opinions, list multiple triplets.
5. Use all words from the original text to answer without any changes.
Example: (generated from training set)
Text: …
Steps:
1. Identify aspects: …
2. Evaluate opinions and sentiments:
3. Summarize:
Please analyze the following text: {sentence}
Table 10: LLM prompts in different methods.
A.10 Detailed Results of GPT Experiments
See Table 11.
Method Combination 14Res 14Lap 15Res 16Res
P R F1 P R F1 P R F1 P R F1
A-O-S 0.5151 0.6519 0.5755 0.3979 0.5009 0.4435 0.4334 0.6309 0.5139 0.5112 0.7101 0.5945
A-O 0.5429 0.6871 0.6066 0.4479 0.5638 0.4992 0.4788 0.6969 0.5676 0.5420 0.7529 0.6303
A-S 0.6234 0.7983 0.7001 0.4955 0.7149 0.5853 0.5539 0.7731 0.6454 0.5853 0.8274 0.6856
GPT-3.5-turbo O-S 0.5790 0.7438 0.6512 0.5016 0.6653 0.5719 0.5228 0.7281 0.6086 0.6020 0.7851 0.6814
few-shots A 0.6758 0.8455 0.7512 0.5805 0.8337 0.6844 0.6220 0.838 0.7140 0.6361 0.8739 0.7363
O 0.6138 0.7938 0.6923 0.5587 0.7505 0.6406 0.5780 0.8048 0.6728 0.6444 0.8404 0.7295
S 0.8251 0.9222 0.8710 0.7907 0.8844 0.8349 0.8286 0.9337 0.878 0.8179 0.9466 0.8776
A-O-S 0.4847 0.5905 0.5324 0.3048 0.4030 0.3471 0.3951 0.5670 0.4657 0.4403 0.6381 0.5210
A-O 0.5178 0.6308 0.5687 0.3566 0.4713 0.4061 0.4368 0.6268 0.5148 0.4711 0.6829 0.5576
A-S 0.5991 0.7807 0.6779 0.4088 0.6199 0.4927 0.5123 0.7685 0.6148 0.5199 0.8075 0.6326
GPT-3.5-turbo O-S 0.5708 0.7143 0.6345 0.4288 0.5768 0.4919 0.4834 0.6711 0.5620 0.5481 0.7511 0.6338
CoT A 0.6594 0.8538 0.7441 0.5142 0.7797 0.6197 0.5901 0.8796 0.7063 0.5702 0.8805 0.6922
O 0.6119 0.7723 0.6828 0.4930 0.6716 0.5686 0.5403 0.7500 0.6281 0.5932 0.8128 0.6858
S 0.7942 0.9374 0.8599 0.7229 0.9046 0.8036 0.7477 0.9222 0.8258 0.7494 0.9585 0.8411
A-O-S 0.4941 0.5915 0.5385 0.3378 0.4233 0.3757 0.3902 0.5608 0.4602 0.4649 0.6693 0.5486
A-O 0.5294 0.6338 0.5769 0.3968 0.4972 0.4413 0.4333 0.6227 0.5110 0.4932 0.7101 0.5821
A-S 0.6257 0.7925 0.6993 0.4518 0.6479 0.5324 0.5306 0.7824 0.6324 0.5451 0.8296 0.6579
GPT-3.5-turbo O-S 0.5735 0.7048 0.6324 0.4681 0.6168 0.5322 0.4936 0.6776 0.5712 0.5769 0.7745 0.6612
CoT+few-shots A 0.6907 0.8691 0.7697 0.5572 0.7991 0.6566 0.6026 0.8704 0.7121 0.5873 0.8783 0.7039
O 0.6186 0.7676 0.6851 0.5319 0.7100 0.6082 0.5479 0.7522 0.6340 0.6149 0.8255 0.7048
S 0.8119 0.9336 0.8685 0.7470 0.8960 0.8147 0.7775 0.9366 0.8497 0.7923 0.9733 0.8735
A-O-S 0.5411 0.6620 0.5955 0.3823 0.4861 0.4280 0.4557 0.6041 0.5195 0.5290 0.7101 0.6063
A-O 0.5757 0.7042 0.6335 0.439 0.5582 0.4915 0.5023 0.6660 0.5727 0.5551 0.7451 0.6362
A-S 0.6777 0.8208 0.7424 0.4962 0.7084 0.5836 0.5872 0.7639 0.6640 0.6019 0.8296 0.6977
GPT-4o O-S 0.6047 0.7532 0.6709 0.4863 0.6337 0.5503 0.5613 0.7325 0.6356 0.6318 0.8106 0.7102
few-shots A 0.7367 0.8679 0.7970 0.5951 0.8445 0.6982 0.6611 0.8218 0.7327 0.6471 0.8761 0.7444
O 0.6389 0.8033 0.7117 0.5517 0.7271 0.6274 0.6185 0.8070 0.7003 0.6667 0.8553 0.7493
S 0.8353 0.9336 0.8817 0.7758 0.8699 0.8202 0.8376 0.9366 0.8844 0.8346 0.9585 0.8923
A-O-S 0.4121 0.5332 0.4649 0.2698 0.3771 0.3146 0.3307 0.5093 0.4010 0.3914 0.5817 0.4679
A-O 0.4331 0.5604 0.4886 0.3122 0.4362 0.3639 0.3614 0.5567 0.4383 0.4162 0.6187 0.4977
A-S 0.6163 0.8278 0.7066 0.4486 0.6976 0.5461 0.5200 0.8125 0.6341 0.5374 0.8429 0.6563
GPT-4o O-S 0.4711 0.6246 0.5371 0.3439 0.4779 0.4000 0.4093 0.5987 0.4862 0.4752 0.6532 0.5502
CoT A 0.6667 0.8703 0.7550 0.5280 0.8143 0.6406 0.5762 0.8750 0.6949 0.5799 0.8916 0.7027
O 0.5004 0.6687 0.5724 0.3951 0.5544 0.4614 0.4558 0.6667 0.5414 0.5077 0.6979 0.5878
S 0.7859 0.9545 0.8620 0.7651 0.9133 0.8327 0.7744 0.9597 0.8571 0.7725 0.9674 0.8590
A-O-S 0.4681 0.5986 0.5254 0.2971 0.4085 0.3440 0.3508 0.5381 0.4247 0.4153 0.6109 0.4945
A-O 0.4965 0.6348 0.5572 0.3454 0.4750 0.4000 0.3911 0.6000 0.4736 0.4437 0.6518 0.5280
A-S 0.6295 0.8314 0.7165 0.4718 0.7235 0.5712 0.5262 0.8125 0.6388 0.5458 0.8429 0.6626
GPT-4o O-S 0.5255 0.6824 0.5937 0.3890 0.5347 0.4504 0.4527 0.6513 0.5342 0.5102 0.6894 0.5864
CoT+few-shots A 0.6795 0.8774 0.7658 0.5467 0.8337 0.6604 0.5898 0.8819 0.7069 0.5858 0.8916 0.7070
O 0.5610 0.7342 0.6360 0.4395 0.6119 0.5116 0.5023 0.7215 0.5923 0.5465 0.7383 0.6281
S 0.7927 0.9431 0.8614 0.7705 0.9220 0.8395 0.7890 0.9481 0.8613 0.7976 0.9703 0.8755
Table 11: Performance of different types of element combinations in ABSA tasks using LLM.
A.11 Case Study
Sentence Ground Truth GPT results Ours Precision Recall
Creamy appetizers–taramasalata, eggplant salad, and Greek yogurt(with cuccumber, dill, and garlic) taste excellent when on warm pitas. [’creamy appetizers’, ’creamy’, ’positive’],
[’creamy appetizers’, ’excellent’, ’positive’],
[’warm pitas’, ’warm’, ’neutral’],
[’taramasalata’, ’creamy’, ’positive’],
[’eggplant salad’, ’excellent’, ’positive’],
[’greek yogurt ( with cuccumber , dill , and garlic )’, ’excellent’,’positive’]
[’appetizers’, ’creamy’, ’neutral’],
[’taramasalata’, ’taste excellent’, ’positive’],
[’eggplant salad’, ’taste excellent’, ’positive’],
[’greek yogurt’, ’taste excellent’, ’positive’],
[’warm pitas’, ’warm’, ’neutral’]
[’creamy appetizers’, ’creamy’, ’positive’],
[’creamy appetizers’, ’excellent’, ’positive’],
[’warm pitas’, ’warm’, ’neutral’],
[’eggplant salad’, ’excellent’, ’positive’],
[’greek yogurt , ’excellent’,’positive’
GPT: 1/5
Ours: 4/5
GPT: 1/6
Ours: 4/6
We left without ever getting service. [’service’, ’without ever’, ’negative’] [’service’, ’without ever getting’, ’negative’] [’service’, ’without ever’, ’negative’] GPT: 0/1
Ours: 1/1
GPT: 0/1
Ours: 1/1
I go out to eat and like my courses, servers are patient and never rush courses or force another drink.’ [’servers’, ’patient’, ’positive’] [’courses’, ’like’, ’positive’],
[’servers’, ’patient’, ’positive’],
[’servers’, ’never rush courses’, ’positive’],
[’servers’, ’never force another drink’, ’positive’]
[’servers’, ’patient’, ’positive’], GPT: 1/5
Ours: 1/1
GPT: 1/1
Ours: 1/1
Table 12: Case study

### v2

@@
 The results demonstrate that the pretraining-finetuning paradigm remains highly effective even in the era of large language models.
+The codebase is available at https://github.com/qiaosun22/MiniConGTS.
+11footnotetext: Corresponding Authors
 1 Introduction
@@
 In the milestone research by Peng et al.
-(2020), the compound ABSA subtasks were consolidated into the Aspect Sentiment Triplet Extraction (ASTE) framework.
+(2020), the compound ABSA subtasks were consolidated into the Aspect Sentiment Triplet Extraction (ASTE) task framework.
 For each input corpus, ASTE outputs triplets in the form (Aspect, Opinion, Polarity), where the Aspect term is the target or entity being discussed, the Opinion term is the sentiment or opinion expressed about the aspect, and Polarity indicates whether the opinion is positive, negative, or neutral.
@@
 ET Xu et al.
-(2020) introduced a position-awared tagging scheme with a conditional random field module, effectively addressing span overlapping issues.
+(2020) introduced a position-aware tagging scheme with a conditional random field module, effectively addressing span overlapping issues.
 Recent joint paradigm methods have refined the ASTE task with the development of proficient Grid Tagging Schemes (GTS).
@@
 An overall description of the training process can be found in Figure 2.
-Basically, our design can be broken down into the Minimalist Grid Tagging Scheme (GTS) and the Token-level Contrastive Learning Strategy.
+Basically our design can be break down into the Minimalist Grid Tagging Scheme (GTS) and the Token-level Contrastive Learning Strategy.
 Tokenize the input sequence S using the Tokenizer $\mathrm{Tk}$ and pass the tokenized sequence through the Pretrained Language Model $\mathrm{PLM}$ (such as BERT) to obtain contextualized representations h:
@@
 Then, the inference phase involves with forming the Minimalist GTS and predicting the corresponding class for each cell.
-Once the GTS is predicted, it can be decoded by the GTS decoder into triplets in natural language form.
-The training phase also introduces a novel contrastive learning strategy, in which similar and dissimilar pairs of contextual representations are distinguished.
-The contrastive loss is then weighted and summed with the tagging loss, which is the classification loss between the predicted and ground-truth tagging schemes.
-Our research benefits from the following two closely intertwined aspects: 1) The use of the Minimalist GTS simplifies the learning process by reducing the number of labels, facilitating faster convergence, and seamlessly gearing the contrastive learning.
+Once the GTS is predicted, it can be decoded by the GTS decoder into the triplets in natural language form.
+The training phase additionally introduces a novel contrastive learning strategy, where similar and dissimilar pairs of contextual representations are distinguished.
+The contrastive loss is then weighted and summed with the tagging loss, which is the classification loss between the predicted and ground truth tagging schemes.
+Our research benefits from the following two closely intertwined aspects: 1) The use of the Minimalist GTS simplifies the learning process by reducing the number of labels, facilitating faster convergence and seamlessly gearing the contrastive learning.
 2) The token-level contrastive learning enhances the model’s ability to distinguish between related and unrelated elements within the input sequence, thereby improving the overall accuracy of the tagging system.
-For a more detailed description of our algorithm pipeline, see the pseudo-code in Appendix A.1.
+For a more detailed description for our algorithm pipeline, see the pseudo code in Appendix A.1.
 Figure 2: An overview of the proposed method, where the “Encoder” denotes for the sequential combination of a Tokenizer and a Pretrained Language Model (PLM).
@@
 Then the each intersection of these marked rows and columns can uniquely represent an identical sentiment triplet.
-Thus, each such triplet can be denoted by a 2-D area (submatirx) in the matrix, where Sentiment Polarity is indicated with POS.
+Thus, each such triplet can be noted by a 2-D area (submatirx) in the matrix, where Sentiment Polarity is indicated with POS.
 (positive), NEU.
@@
 $tag_{t}=\hat{\textit{{tag}}}_{i,j;\textit{{tag}}_{i,j}}$ (4)
-3.3 Contrstive Learning Strategy
+3.3 Contrastive Learning Strategy
 3.3.1 Contrastive Learning Label Matrix
@@
 GPT-3.5-turbo zero-shot 44.88 55.13 49.48 30.04 41.04 34.69 36.02 53.40 43.02 39.92 57.78 47.22
-GPT-3.5-turbo few-shots 51.51 65.19 57.55 39.79 50.09 44.35 43.34 63.09 51.39 51.12 71.01 59.45
+GPT-3.5-turbo few-shot 51.51 65.19 57.55 39.79 50.09 44.35 43.34 63.09 51.39 51.12 71.01 59.45
 GPT-3.5-turbo CoT 48.47 59.05 53.24 30.48 40.30 34.71 39.51 56.70 46.57 44.03 63.81 52.10
-GPT-3.5-turbo CoT+few-shots 49.41 59.15 53.85 33.78 42.33 37.57 39.02 56.08 46.02 46.49 66.93 54.86
+GPT-3.5-turbo CoT+few-shot 49.41 59.15 53.85 33.78 42.33 37.57 39.02 56.08 46.02 46.49 66.93 54.86
 GPT-4o zero-shot 32.99 38.13 35.37 17.81 22.55 19.90 27.85 37.73 32.05 32.17 43.00 36.80
-GPT-4o few-shots 54.11 66.20 59.55 38.23 48.61 42.80 45.57 60.41 51.95 52.90 71.01 60.63
+GPT-4o few-shot 54.11 66.20 59.55 38.23 48.61 42.80 45.57 60.41 51.95 52.90 71.01 60.63
 GPT-4o CoT 41.21 53.32 46.49 26.98 37.71 31.46 33.07 50.93 40.10 39.14 58.17 46.79
-GPT-4o CoT+few-shots 46.81 59.86 52.54 29.71 40.85 34.40 35.08 53.81 42.47 41.53 61.09 49.45
+GPT-4o CoT+few-shot 46.81 59.86 52.54 29.71 40.85 34.40 35.08 53.81 42.47 41.53 61.09 49.45
 Ours
@@
 We evaluate ASTE performance using the widely accepted (Precision, Recall, F1) metrics.
-The results of the data set $\mathcal{D}_{2}$ are shown in Table 1, while the results of $\mathcal{D}_{1}$ are presented in Appendix Table 8.
+Results of the dataset $\mathcal{D}_{2}$ are shown in Table 1, while the results of $\mathcal{D}_{1}$ are presented in Appendix Table 8.
 The best results are highlighted in bold, and the second-best results are underscored.
 Our proposed method consistently achieves state-of-the-art performance or ranks second in most evaluated cases.
-In particular, on the data set $\mathcal{D}_{1}$, the proposed method achieves a substantial improvement of 3.
-08% in the F1 score in the 14Lap subset.
+Notably, on dataset $\mathcal{D}_{1}$, the proposed method achieves a substantial 3.08% improvement in F1 score on the 14Lap subset.
 This improvement is particularly significant given that the highest score on this subset is the lowest among all datasets, showcasing our method’s effectiveness in handling challenging instances.
 Moreover, on the 14Res subset, our F1 score exceeds 76.00, which, to the best of our knowledge, is the highest reported performance.
-For the dataset $\mathcal{D}_{2}$, our method outperforms all state-of-the-art approaches by more than 1 percentage point on the 14Res, 14Lap, and 16Res subsets.
+For dataset $\mathcal{D}_{2}$, our method outperforms all state-of-the-art approaches by over 1 percentage point on the 14Res, 14Lap, and 16Res subsets.
 Only on the 16Res subset does the BDTF method (Zhang et al., 2022b) achieve a slightly better performance.
@@
 Our approach not only requires less memory usage for higher performance compared to traditional ASTE methods but also offers much faster runtime even using a relatively lower-cost GPU.
-Table 3 provides another comparative analysis of the tagging schemes.
+Table 3 provides another comparative analysis of tagging schemes.
 Our method has a compact design with the fewest classes of labels.
-Moreover, compared with other SOTA baselines, our method does not rely on any additional linguistic information enhancement.
+What’s more, compared with other SOTA baselines, our method does not rely on any additional linguistic information enhancement.
 5.4 Case Study
 A case analysis is presented in Table 12, where the proposed method demonstrates solid performance. Despite minor faults in missing the full terms, it exhibits a profound understanding of the case.
-It is quite interesting to investigate the GPT error cases.
+It is quite interesting to investigate the GPT’s error cases.
 The findings reveal that the performance of the GPT model is mixed - while it is able to identify more aspect-opinion (A-O) pairs than the ground truth annotations in some cases, this comes at the cost of a decreased precision.
@@
 Furthermore, the GPT model appears to be overly sensitive to the presence of adverbs (such as “very”, “a bit”, etc.) in the input.
-This sensitivity manifests itself in the model frequently adding or removing adverbs when extracting the opinion components, which further contributes to a decrease in the overall accuracy of the ABSA task.
+This sensitivity manifests in the model frequently adding or removing adverbs when extracting the Opinion components, which further contributes to a decrease in the overall accuracy of the ABSA task.
 These findings highlight the importance of developing ABSA models that can strike the right balance between extracting all relevant aspect-opinion pairs, while still maintaining a high degree of precision.
@@
 6 Conclusion
-In this work, we have introduced an elegant and efficient framework for ASTE to achieve SOTA performance.
+In this work, we have introduced an elegant and efficient framework for ASTE, achieving SOTA performance.
 Our approach is built upon two effective components: a new tagging scheme and a novel token-level contrastive learning implementation.
@@
 7 Limitations
-Our method is based on a 2D matrix tagging scheme, where the time complexity for decoding, given the input corpus length $N$, is $O(N^{2})$.
+Our method is based on a 2D-matrix tagging scheme, where the time complexity for decoding, given the input corpus length $N$, is $O(N^{2})$.
 This may be unacceptable when $N$ is too large.
@@
 We reserve the right to limit or modify the use of our technology to prevent misuse.
+9 Acknowledgement
+This work was partially supported by the National Key R&D Program of China under Grant No.
+2022ZD0160101.
+The authors would also like to thank Shanghai Artificial Intelligence Laboratory for their support and resources, as well as colleagues from Shanghai Jiao Tong University for valuable discussions.
+This work was conducted during the internships of Qiao Sun, Liujia Yang, and Minghao Ma at the Shanghai Artificial Intelligence Laboratory.
 Appendix A Appendix
@@
 2:  Tensor Operations: ||×$\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}=\texttt{expand}(\mathcal{H}_{|\mathcal{H}|})$, ||×$\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T}=\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}.\texttt{transpose}()$;
-3:  Similarity matrix:
-$\mathbf{Sim}_{|\mathcal{H}|\times|\mathcal{H}|}=-(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})\circ(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})$
-where
+3:  Similarity matrix: $\mathbf{Sim}_{|\mathcal{H}|\times|\mathcal{H}|}=$ (||×||||×$-(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})\circ(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})$ where
 $\mathbf{Sim}_{i,j}=-\left\lVert\mathcal{H}_{i}-\mathcal{H}_{j}\right\rVert^{2}$
@@
 Lemma 2.
-In a 2D-tagging for ASTE, at least three basic goals must be met: 1) correctly identifying the (Aspect, Opinion) pairs, 2) correctly classifying the sentiment polarity of the pair based on the context, and 3) avoiding boundary errors, such as overlapping [^1] It occurs when one single word belongs to multiple classes in different triplets.
-, confusion [^2] It occurs when there is a lack of location restrictions so that multiple neighbored candidates can not be uniquely distinguished.
-, and conflict [^3] It occurs when one single word is composed of multiple tokens, and the predict gives predictions that are not aligned with the word span.
+In a 2D-tagging for ASTE, at least three basic goals must be met: 1) correctly identifying the (Aspect, Opinion) pairs, 2) correctly classifying the sentiment polarity of the pair based on the context, and 3) avoiding boundary errors, such as overlapping***It occurs when one single word belongs to multiple classes in different triplets.
+, confusionIt occurs when there is a lack of location restrictions so that multiple neighbored candidates can not be uniquely distinguished.
+, and conflictIt occurs when one single word is composed of multiple tokens, and the predict gives predictions that are not aligned with the word span.
 .
@@
 This tagging is crucial to both identify the beginning of an triplet and label the sentiment polarity.
-Figure 6 shows a comprehensive case of our tagging scheme, in which the left matrix is an appearance of our tagging scheme and can be decomposed into two separate components.
+Figure 6 shows a comprehensive case of our tagging scheme, in which the left matrix is an appearance of our tagging scheme, and it can be decomposed into two separate components.
 The middle matrix is the first component, which takes only one tag to locate the up-left beginning of an area, and the second component simply predicts a binary classification to figure out the full area.
@@
     Assuming Incorrect Identification: Assume that some (Aspect, Opinion) pairs are incorrectly identified in $M$.
-This would mean that there exists at least one pair $(i,j)$ where $M[i][j]$ does not represent the actual relationship (Aspect, Opinion) in $S$.
+This would mean that there exists at least one pair $(i,j)$ where $M[i][j]$ does not represent the actual (Aspect, Opinion) relationship in $S$.
 This misrepresentation leads to incorrect sentiment analysis results, which contradicts the requirement of the task to provide accurate sentiment analysis, thereby proving that our identification must be correct.
@@
 Conclusion
-The contraposition approach solidifies that the ASTE tagging strategy in a 2D labeling framework successfully achieves the correct identification of pairs, accurate classification of sentiment, and effective management of boundary errors, as any failure in these aspects leads to contradictions with the task requirements.
+The contraposition approach solidifies that the tagging strategy for ASTE in a 2D labeling framework successfully achieves the correct identification of pairs, accurate classification of sentiment, and effective management of boundary errors, as any failure in these aspects leads to contradictions with the task requirements.
 A.6 Baselines
@@
 The provided text: {sentence}"
-few-shots Building upon the zero-shot method, a small number of examples from the training set are added to the prompted sentence: "Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment].
+few-shot Building upon the zero-shot method, a small number of examples from the training set are added to the prompted sentence: "Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment].
 For example: input: {train sentence} output: {train triplets}, … (some other examples).
 You only need to provide the triplets, no additional explanations are required.
-The provided text: {sentence}"
+The provided text: {sentence}".
+We utilized 5-shot, 10-shot, and 20-shot methods, all randomly sampled from the training set.
+The results indicate that the 5-shot method performed the best, while the performances of the 10-shot and 20-shot methods showed a decline.
+The tables presents the output results for the 5-shot method.
 Table 6: Baseline methods with brief introduction.
@@
 GPT-3.5-turbo zero-shot 39.21 56.17 46.18 26.21 40.69 31.88 31.21 52.75 39.21 35.28 59.64 44.34
-GPT-3.5-turbo few-shots 50.32 64.75 56.63 29.67 43.90 35.41 36.94 61.01 46.02 44.80 69.96 54.62
+GPT-3.5-turbo few-shot 50.32 64.75 56.63 29.67 43.90 35.41 36.94 61.01 46.02 44.80 69.96 54.62
 GPT-3.5-turbo CoT 40.78 57.93 47.86 28.37 43.25 34.27 35.17 57.11 43.53 40.32 65.79 50.00
-GPT-3.5-turbo CoT+few-shots 44.97 57.81 50.59 28.31 43.04 34.15 35.71 58.49 44.35 43.72 66.45 52.74
+GPT-3.5-turbo CoT+few-shot 44.97 57.81 50.59 28.31 43.04 34.15 35.71 58.49 44.35 43.72 66.45 52.74
 Ours
@@
 Sentence Ground Truth Predictions
-GPT-3.5 zero-shot GPT-3.5 few-shots GPT-4 zero-shot GPT-4 few-shots
+GPT-3.5 zero-shot GPT-3.5 few-shot GPT-4 zero-shot GPT-4 few-shot
 It is a cozy place to go with a couple of friends. [place, cozy, positive] [place, cozy, positive],
@@
 The provided text: {sentence}
-few-shots Suppose you are an expert of aspect-based sentiment analysis. Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. For example:
+few-shot Suppose you are an expert of aspect-based sentiment analysis. Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. For example:
 input: The food is uniformly exceptional , with a very capable kitchen which will proudly whip up whatever you feel like eating , whether it ’s on the menu or not .
@@
 Please analyze the following text: {sentence}
-CoT+few-shots Suppose you are an expert of aspect-based sentiment analysis. Please analyze the given text for aspect-based sentiment analysis using the following steps:
+CoT+few-shot Suppose you are an expert of aspect-based sentiment analysis. Please analyze the given text for aspect-based sentiment analysis using the following steps:
 Definitions:
@@
 GPT-3.5-turbo O-S 0.5790 0.7438 0.6512 0.5016 0.6653 0.5719 0.5228 0.7281 0.6086 0.6020 0.7851 0.6814
-few-shots A 0.6758 0.8455 0.7512 0.5805 0.8337 0.6844 0.6220 0.838 0.7140 0.6361 0.8739 0.7363
+few-shot A 0.6758 0.8455 0.7512 0.5805 0.8337 0.6844 0.6220 0.838 0.7140 0.6361 0.8739 0.7363
 O 0.6138 0.7938 0.6923 0.5587 0.7505 0.6406 0.5780 0.8048 0.6728 0.6444 0.8404 0.7295
@@
 GPT-3.5-turbo O-S 0.5735 0.7048 0.6324 0.4681 0.6168 0.5322 0.4936 0.6776 0.5712 0.5769 0.7745 0.6612
-CoT+few-shots A 0.6907 0.8691 0.7697 0.5572 0.7991 0.6566 0.6026 0.8704 0.7121 0.5873 0.8783 0.7039
+CoT+few-shot A 0.6907 0.8691 0.7697 0.5572 0.7991 0.6566 0.6026 0.8704 0.7121 0.5873 0.8783 0.7039
 O 0.6186 0.7676 0.6851 0.5319 0.7100 0.6082 0.5479 0.7522 0.6340 0.6149 0.8255 0.7048
@@
 GPT-4o O-S 0.6047 0.7532 0.6709 0.4863 0.6337 0.5503 0.5613 0.7325 0.6356 0.6318 0.8106 0.7102
-few-shots A 0.7367 0.8679 0.7970 0.5951 0.8445 0.6982 0.6611 0.8218 0.7327 0.6471 0.8761 0.7444
+few-shot A 0.7367 0.8679 0.7970 0.5951 0.8445 0.6982 0.6611 0.8218 0.7327 0.6471 0.8761 0.7444
 O 0.6389 0.8033 0.7117 0.5517 0.7271 0.6274 0.6185 0.8070 0.7003 0.6667 0.8553 0.7493
@@
 GPT-4o O-S 0.5255 0.6824 0.5937 0.3890 0.5347 0.4504 0.4527 0.6513 0.5342 0.5102 0.6894 0.5864
-CoT+few-shots A 0.6795 0.8774 0.7658 0.5467 0.8337 0.6604 0.5898 0.8819 0.7069 0.5858 0.8916 0.7070
+CoT+few-shot A 0.6795 0.8774 0.7658 0.5467 0.8337 0.6604 0.5898 0.8819 0.7069 0.5858 0.8916 0.7070
 O 0.5610 0.7342 0.6360 0.4395 0.6119 0.5116 0.5023 0.7215 0.5923 0.5465 0.7383 0.6281
@@
 A.11 Case Study
+See Table 12.
 Sentence Ground Truth GPT results Ours Precision Recall